A company estimates that 30% of the country has seen its commercial and that if a person
is the probability that a person chosen at random in the country will have seen the
sees its commercial, there is a 20% probability that the person willbuy its product. What
commercial and bought the product?
1
Expert's answer
2021-09-14T06:09:23-0400

"P(buys|seen) = 0.2 \\

P(seen) = 0.3 \\

P(A \cap B) = P(B) \times P(A|B)"

For independent events

"P(A \cap B) = P(B) \times P(A) \\

P(buys \cap seen) = P(seen) \times P(buys|seen) \\

= 0.3 \times 0.2 \\

= 0.06"